A monopolistic producer of two goods, 1 and 2, has a joint total cost function




where and denote the quantity of items of goods 1 and 2, respectively that are produced. If P1 and P2 denote the corresponding prices then the demand equations are




Using the Lagrange multiplier approach, find the maximum profit if the firm is contracted to produce a total of 15 goods of either type. Estimate the new optimal profit if the production quota rises by 1 unit.


1
Expert's answer
2021-05-25T10:39:40-0400

"TC=10Q_1+Q_1Q_2+10Q_2"


Then "max\pi ~~~~~~~ Q_1+Q_2\le 15"


Then "L=P_1Q_1+P_2Q_2-TC +\lambda [15-Q_1-Q_2]"


"L=40Q_1-Q_1^2-Q_2^2+30Q_2-10Q_2+2Q_1Q_2 +\lambda [15-Q_1-Q_2]
\\[9pt]


L=40Q_1-Q_1^2-Q_2^2+20Q_2+2Q_1Q_2+\lambda[15-Q_1-Q_2]"



Now,

"\dfrac{dL}{dQ_1}=40-2Q_1+2Q_2-\lambda=0"


and "\dfrac{dL}{dQ_2}=-2Q_2+20+2Q_1-\lambda=0"


So From FOC:


"40-2Q_1+2Q_2=20+2Q_1-2Q_2

\\[9pt]

\Rightarrow 20=4Q_1-4Q_2

\\[9pt]

\Rightarrow Q_1-Q_2=5"


and Q_1+Q_2=15


Solving above equation we get-


"Q_1=10,Q_2=5"


Max "\pi =P_!Q_1+P_2Q_2-TC"


"P_1=50-10+5=45, P_2=30+20-5=45"


Then,

"\pi=45(15)-10(10)-50-50
\\
 =675-100-100=475"


Now "\lambda=40-2(Q_1-Q_2)"

      "=40-2(5)\n\n =30"


So if quota is incresed by 1 unit, then optimal output is increased by $30.